# A Generalization of Lévy’s Theorem on Transition Matrices

### 07 Jan 2024 v1

@@ Abstract.
We generalize a fundamental theorem on transition matrices stating that each component is either strictly positive for all times or identically zero (”Lévy’s Theorem“). Our proof of this fact that does not require the matrices to be Markovian nor to be continuous at time zero. We also provide a formulation of this theorem in the terminology of one-parameter operator semigroups.
@@
transition matrix; operator semigroup; discrete spaces
@@ 1. Introduction
Let $I$ denote a countable set. A family of functions $p_{ij}\colon(0,\infty)\to[0,\infty)$, $i,j\in I$, is called a transition matrix (over the index set $I$) if for every $i,j\in I$ the following two conditions are fulfilled:
@@
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Then for all $i,j\in I$
@@
This statement can directly be transferred to the theory of operator semigroups as follows (see Corollary 3.1 and Example 3.2):
@@
Let $E$ be one of the sequence spaces $c_{0}$, $c$ or $\ell^{p}$ for some $1\leq p\leq\infty$ (or, more precisely, a discrete Archimedean Riesz in the sense of Corollary 3.1). For $k\in\mathds{N}$ denote by $e_{k}$ the $k$-th canonical unit vectors of $E$. Let $(T_{t})_{t>0}$ be a one-parameter semigroup of positive linear operators on $E$ such that $t\mapsto\langle T_{t}e_{i},e_{j}\rangle$ is continuous on $(0,\infty)$ for all $i,j\in\mathds{N}$ (this is obviously fulfilled whenever the semigroup is strongly continuous).
Then, for every $i,j\in\mathds{N}$, either $\langle T_{t}e_{i},e_{j}\rangle>0$ for all $t>0$ or $\langle T_{t}e_{i},e_{j}\rangle=0$ for all $t>0$.
Under additional assumptions, Theorem 1.1 is well-known in the literature. A first partial proof was given by Lévy in [14], which is why Theorem 1.1 is sometimes referred to as ”Levy’s Theorem“. In [6, Thm 10.1], Chung proved Statement (1.1) in the most general situation so far, where he assumed the transition matrix in addition to be continuous at $t=0$, i.e.
@@
    $\lim_{t\to 0}p_{ij}(t)=\begin{cases}0&i\neq j\\ 1&i=j\end{cases}$,
Such transition matrices satisfying property (c) are also called standard in the literature. Unfortunately, the proof given in [6, Thm 10.1] is apparently not complete: it is unclear how to conclude $p(2t_{0})\to 0$ as $N\to 0$ from the uniform convergence of (10.2) at the top of page 76, as several objects depend on $N$; however, it is not difficult to understand the author’s intention by taking his previous publications as [7, Thm II.1.5] into account. We also refer to the notes in [7] for further historical references and several different proofs of Theorem 1.1 for transition matrices which are Markovian, i.e.
@@
    $\sum_{j\in I}p_{ij}(t)=1$ for all $t>0$.
In the theory of operator semigroups a version of Theorem 1.2 for positive analytic $C_{0}$-semigroups on general (not necessarily discrete) Banach lattices can be found in [2, Thm C-III 3.2(b)]. For self-adjoint $C_{0}$-semigroups on $L^{2}$-spaces this result is due to Simon [16, Thm 1]. For $C_{0}$-semigroups on $L^{p}$-spaces it was shown by Kishimoto and Robinson [11, Lem 8] and it was generalized to the Banach lattice setting by Majewski and Robinson [15, Prop 2]. Jochen Glück pointed out to the author that, similar to the main result of the present article, the assumption of strong continuity at time zero can also be dropped in the analytic case. This result is provided as Theorem A.1 in the appendix. In the 21th century, results similar to Theorem 1.2 have also been proven in the context of infinite graphs, cf. [9, Thm 7.3] and [10, Thm 1.26], which is a special case of both, the discrete and the analytic setting.
The present article has three goals: First, to give a complete proof of Chung’s result. Second, to generalize it by dropping both assumptions (c) and (d). And third, to transfer the result to the theory of operator semigroups, where it is apparently little known.
2. Transition Matrices
The proof of the following theorem is a combination and adaption of [7, Thm II.1.5] and [6, Thm 10.1].
@@ 2. Transition Matrices
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$. Then $p_{ij}(t)=0$ for all $0<t\leq t_{0}$ implies $p_{ij}(t)=0$ for all $t>0$.
@@
Let $N\coloneqq\#F$ denote the number of elements in $F$. By enlarging the set $F$ if necessary we may assume that $N$ is an even number. We define the stepwidth $s\coloneqq\frac{t_{0}}{4N}$ and the abbreviations $N_{0}\coloneqq 8\delta N$ and $N_{1}\coloneqq 8(1-\delta)N$. Note that $N_{0},N_{1}\in\mathds{N}$, $N_{0}s=\delta 2t_{0}$, $N_{1}s=(1-\delta)2t_{0}$ and $(N_{0}+N_{1})s=2t_{0}$.
@@
By definition of $C_{m}$, $u(m+1,n)\leq u(m,n)$ for all $m,n$ and $u(m,n)=0$ whenever $n\leq m$. Moreover, $v(m,n)<\varepsilon$ whenever $D_{m}$ is disjoint from $F$ and $n\in\{N_{0},\dots,N_{1}\}$. Given $M,m\in\mathds{N}$, $k\in C_{M}$ and $l\in I$ such that $p_{lk}(ms)>0$, then $l\in C_{M-m}$. In fact, otherwise it would follow from
$p_{ik}(M^{\prime}s)\geq p_{il}\big{(}(M^{\prime}-m)s\big{)}p_{lk}(ms)>0\text{ for some }m+1\leq M^{\prime}\leq M$
that $k\not\in C_{M}$. From this, we now conclude the following two assertions: First,
@@
Recall that at most $N$ of the disjoint sets $D_{m}$ can have a non-empty intersection with $F$. As the set $\{N_{0},\dots,4N-N_{0}\}$ of all choice for $m$ has $4N-2N_{0}+1\geq(4-16\delta)N=3N$ elements, we find $M\subseteq\{N_{0},\dots,4N-N_{0}\}$ such that #$\#M=N$ and $D_{m}\subseteq I\setminus F$ for all $m\in M$. Summing over $m\in M$, we obtain from (2.1)
$Nc$ $\leq\sum_{m\in M}\sum_{n=N_{0}+1}^{N_{1}}v(m,n)=\sum_{n=N_{0}+1}^{N_{1}}\sum_{m\in M}v(m,n)$
@@
$<\sum_{n=N_{0}+1}^{N_{1}}\varepsilon<(N_{1}-N_{0})\varepsilon=(8N-16\delta N)\varepsilon=7N\varepsilon<Nc,$
@@
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t_{0})=0$. Then $p_{ij}(t)=0$ for all $t\geq t_{0}$
@@
In order to obtain Theorem 1.1, it remains to show that $p_{ij}(t_{1})>0$ implies that $p_{ij}(t)>0$ for all $t\geq t_{1}$; a fact that is easy to see for Markovian (or standard) transition matrices, cf. [7, Thm II.1.4.], but surprisingly sophisticated to achieve in the general setting of this article. This is the aim of what follows.
@@
Let $p_{ij}$ be a transition matrix over a countable index set $I$ and let $j\in I$ and assume that $p_{jj}(t)=0$ for all $t>0$.
@@
The assumption $p_{ij}(t_{1})>0$ for some $t_{1}>0$ in Part (b) if Lemma 2.3 may look a bit artifical at first glance. However, the example
@@
of a constant transition matrix shows that this assumption is indeed necessary. The next theorem contains the article’s main result.
@@
Let $p_{ij}$ be a transition matrix over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t_{0})=0$. Then $p_{ij}(t)=0$ for all $t>0$.
@@
and $p_{kj}(u-s)>0$ that $p_{kk}(s)=0$. Now Theorem 2.1 implies that $p_{kk}(t)=0$ for all $t>0$. Theorefore, $p_{kl}(t)=0$ for all $t>0$ und $l\in I$ by Part (b) of Lemma 2.3. In paticular, $p_{kj}$ is identically zero in contradiction to our assumption.
@@
3. Operator Semigroups on Discrete Spaces
@@
In what follows, we give a reformulation of Theorem 2.4 for one-parameter semigroups on discrete Riesz spaces. Recall that a family of linear operators $(T_{t})_{t\in(0,\infty)}$ on a vector space $X$ is called a (one-parameter operator) semigroup if $T_{t+s}=T_{t}T_{s}$ for all $t,s>0$.
@@
One-parameter operator semigroups appear naturally to describe solutions of linear autonomous evolution equations. In many applications the underlying Banach space is a function space and thus exhibits some kind of order structure. If, in such a situation, a positive initial value of the evolution equation leads to a positive solution, one speaks of a positive operator semigroup. We refer to the recent monograph [5] for a modern introduction to the theory of positive semigroups.
We recall some basic notations from the theory of Riesz spaces according to [13]. A real ordered vector space is called a Riesz space (or vector lattice) if for every pair $u,v\in E$ the supremum $u\vee v\coloneqq\sup\{u,v\}$ exists in $E$. In this case, one defines $\lvert u\rvert\coloneqq u\vee(-u)$ and $u\wedge v\coloneqq-\big{(}(-u)\vee(-v)\big{)}$ for all $u,v\in E$ and $E_{+}\coloneqq\{u\in E:u\geq 0\}$. Two vectors $u,v\in E$ are called disjoint if $\lvert u\rvert\wedge\lvert v\rvert=0$. A linear operator $T\colon E\to F$ between Riesz spaces $E$ and $F$ is called positive if $TE_{+}\subseteq F_{+}$; an operator semigroup is called positive if it consists of positive operators. A Riesz space $E$ is called Archimedean if for every $u\in E_{+}$ such that the set $\{nu:n\in\mathds{N}\}$ is order bounded in $E$ it follows that $u=0$.
Let $E$ be an Archimedean Riesz space. A vector subspace $J\subseteq E$ such that $\lvert u\rvert\leq\lvert v\rvert$ for $v\in J$ implies that $u\in J$ is called an ideal. The smallest ideal containing a given $u\in E_{+}$ is called the principal ideal generated by $u$ and is given by $E_{u}=\{v\in E:\exists c\geq 0\text{, }\lvert v\rvert\leq cu\}$. If an ideal $J\subset E$ is closed under general suprema, i.e., $\sup A\in J$ for each set $A\subseteq J$ whose supremum exists in $E$, then $J$ is called a band. For each subset $A\subseteq E$ we denote by $A^{\bot}\coloneqq\{v\in E:\lvert v\rvert\wedge\lvert u\rvert=0\text{ for all }u\in A\}$ the disjoint complement of $A$; we recall that $A^{\bot}$ is a band and that $A^{\bot\bot}\coloneqq(A^{\bot})^{\bot}$ is the smallest band containing $A$. A vector $u\in E_{+}$ is called a weak (order) unit (of $E_{+}$) if the principal band $\{u\}^{\bot\bot}$ is equal to the whole space $E$.
An element $d\in E$, $d\neq 0$, is called discrete if the principal ideal $E_{d}$ is one-dimensional, i.e. $E_{d}=\{\lambda d:\lambda\in\mathds{R}\}$. Since $E$ is Archimedean, the principal band generated by a discrete element $d$ is given by $\{d\}^{\bot\bot}=\{\lambda d:\lambda\in\mathds{R}\}$ and is a projection band, i.e. $\{d\}^{\bot\bot}$ is the range of a positive projection $P_{d}\colon E\to\{d\}^{\bot\bot}$ such that $I-P_{d}$ is a positive projection onto $\{d\}^{\bot}$ [13, Thm 26.4]. Since the band projection is uniquely determined, for every discrete element $d\in E$ there exists a uniquely determined linear functional $\varphi_{d}\colon E\to\mathds{R}$ such that $\langle\varphi_{d},u\rangle d=P_{d}u$ for all $u\in E$. Clearly, $\varphi_{d}$ is positive if and only if $d\in E_{+}$. We call the Archimedean Riesz space $E$ discrete if there exists a set $D\subseteq E$ of discrete elements such that $D^{\bot\bot}=E$.
Using this terminology, Theorem 2.4 can be reformulated as follows.
@@
Let $E$ be a discrete Archimedean Riesz space and $D\subseteq E_{+}$ be a countable set of pairwise disjoint discrete elements such that $D^{\bot\bot}=E$. Let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup on $E$ such that $t\mapsto\langle\varphi_{d},T_{t}d^{\prime}\rangle$ is continuous on $(0,\infty)$ for all $d,d^{\prime}\in D$.
Then for all $d\in D$, either $T_{t}d$ is weak order unit for all $t>0$ or $T_{t}d=0$ for all $t>0$.
Example 3.2.
Let $E$ denote one of the following sequence spaces: $c$ (space of convergent sequences), $c_{0}$ (the space of nullsequences), $l^{p}$ (space of $p$-summable sequences for $1\leq p<\infty$) or $\ell^{\infty}$ (space of bounded sequences). Let $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup that is strongly continuous, i.e. $t\mapsto{T_{t}x}$ is continuous on $(0,\infty)$ for all $x\in E$. Then for all $x\in E_{+}$, either $T_{t}x$ is a weak order unit for all $t>0$ or $T_{t}x=0$ for all $t>0$.
@@
While Corollary 3.1 follows immediately from Theorem 2.4 applied to the transition matrices $p_{ij}(t)=\langle T_{t}d_{i},\varphi_{d_{j}}\rangle$ (for discrete elements $d_{i},d_{j}\in D$), it is also possible to conclude Theorem 2.4 from the Corollary as follows: Let $p_{ij}$ be a transition matrix over the index set $\mathds{N}$. On $\ell^{0}$, the discrete Archimedean Riesz space of all real sequences, we consider the set
$E\coloneqq\{(x_{n})\in\ell^{0}:\sum\nolimits_{n}\lvert x_{n}\rvert p_{nm}(t)<\infty\text{ for all }t>0\text{ and }m\in\mathds{N}\},$

### 17 Mar 2024 v2

# A Generalization of Lévy’s Theorem on positive matrix semigroups

@@ Abstract.
We generalize a fundamental theorem on positive matrix semigroups stating that each component is either strictly positive for all times or identically zero (“Lévy’s Theorem”). Our proof of this fact that does not require the matrices to be continuous at time zero. We also provide a formulation of this theorem in the terminology of positive operator semigroups on sequence spaces.
@@
transition matrix; matrix semigroup; operator semigroup; discrete space; sequence space
@@ 1. Introduction
Let $I$ denote a countable set. A family of functions $p_{ij}\colon(0,\infty)\to[0,\infty)$, $i,j\in I$, is called a positive matrix semigroup (or transition semigroup) over the index set $I$ if for every $i,j\in I$ the following two conditions are fulfilled:
@@
Let $p_{ij}$ be a positive matrix semigroup over a countable index set $I$. Then for all $i,j\in I$
@@
We also transfer this theorem to the theory of operator semigroups, where it becomes the following (see Corollary 3.1):
@@
Let $E$ be a sequence space (as defined in Section 3). Let $(T_{t})_{t\in(0,\infty)}$ be a positive one-parameter semigroup of sequentially order continuous operators on $E$ such that $t\mapsto(T_{t}e^{i})_{j}$ is continuous on $(0,\infty)$ for all $i,j\in\mathds{N}$.
Then for all $i,j\in\mathds{N}$ either $(T_{t}e^{i})_{j}>0$ for all $t>0$ or $(T_{t}e^{i})_{j}=0$ for all $t>0$.
Here, $e^{i}\coloneqq\big{(}\mathds{1}_{\{i\}}(j)\big{)}_{j\in\mathds{N}}$ denotes the $i$-th unit vector. On sequence spaces with a complete and order continuous lattice norm, such as $c_{0}$ (space of null sequences) or $\ell^{p}$ (space of $p$-summable sequences for $1\leq p<\infty$), the result simplifies as follows (see Corollary 3.3):
Corollary 1.3.
Let $E$ denote a sequence space with a complete and order continuous lattice norm and let $(T_{t})_{t\in(0,\infty)}$ be a positive one-parameter semigroup such that $t\mapsto(T_{t}e^{i})_{j}$ is measurable $(0,\infty)$ for all $i,j\in\mathds{N}$. Then, for every $i,j\in\mathds{N}$, either $(T_{t}e^{i})_{j}>0$ for all $t>0$ or $(T_{t}e^{i})_{j}=0$ for all $t>0$.
A version for spaces with uncountably many atoms, such as $\ell^{p}(\mathds{R})$ for $1\leq p<\infty$, can be found as Corollary 3.5.
Under additional assumptions, Theorem 1.1 is well-known in the literature. A first partial proof was given by Lévy in [16, Thm II.8.1] for positive matrix semigroups which are Markovian, i.e.
@@
    $\sum_{j\in I}p_{ij}(t)=1$ for all $t>0$ and $i\in I$,
and the additional assumption that all $p_{jj}$ have finite derivatives at time 00. This is why Theorem 1.1 is sometimes referred to as “Levy’s Theorem”. After Chung had pointed out some inaccuracies to him, Lévy amended his arguments for the general Markovian case in [17, Sec 7].
According to the notes in [6], an alternative proof for Markovian matrix semigroups was found independently by D. G. Austin, simplified by Chung using probabilistic methods and published as [6, Thm II.5.2]. Another algebraic proof was found by D. Ornstein and published as [6, Thm II.1.5]. Both proofs make use of the fact that for a Markovian matrix semigroup for all $i,j$ the limits $\lim_{t\downarrow 0}p_{ij}(t)$ exist, cf. [6, Thm II.1.3].
In [7, Thm 10.1], Chung proved Statement (1.1) in the most general situation so far, where the Markov condition (c) is replaced by continuity at $t=0$, i.e.
@@
    $\lim_{t\to 0}p_{ij}(t)=\begin{cases}0&i\neq j\\ 1&i=j\end{cases}$ for all $i,j\in I$.
Matrix semigroups satisfying property (d) are also called standard in the literature. Unfortunately, the proof given in [7, Thm 10.1] is apparently not complete: it is unclear how to conclude $p(2t_{0})\to 0$ as $N\to 0$ from the uniform convergence of (10.2) at the top of page 76, as several objects depend on $N$. This has recently led to some uncertainty regarding correctness of the theorem within the operator semigroup community. However, it is not difficult to understand the author’s intention by taking his previous publications as [8, Thm II.1.5] into account.
In the theory of operator semigroups a version of Theorem 1.2 for positive analytic $C_{0}$-semigroups on general (not necessarily discrete) Banach lattices can be found in [2, Thm C-III 3.2(b)]. For self-adjoint $C_{0}$-semigroups on $L^{2}$-spaces this result is due to Simon [20, Thm 1]. For $C_{0}$-semigroups on $L^{p}$-spaces it was shown by Kishimoto and Robinson [13, Lem 8] and it was generalized to the Banach lattice setting by Majewski and Robinson [18, Prop 2]. Jochen Glück pointed out to the author that, similar to the main result of the present article, the assumption of strong continuity at time zero can also be dropped in the analytic case. This result is provided as Theorem A.1 in the appendix.
In the 21th century, results similar to Theorem 1.2 have also been proven in the context of infinite graphs, cf. [10, Thm 7.3] and [12, Thm 1.26], which is a special case of both, the discrete and the analytic setting.
Summarizing, the present article has three goals: First, to give a complete proof of Chung’s result. Second, to generalize it by dropping both assumptions (c) and (d). And third, to transfer the theorem to the theory of operator semigroups, where it is apparently little known.
2. Positive Matrix Semigroups
The proof of the following theorem is a combination and adaption of [8, Thm II.1.5] and [7, Thm 10.1].
@@ 2. Positive Matrix Semigroups
Let $p_{ij}$ be a positive matrix semigroup over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t)=0$ for all $0<t\leq t_{0}$. Then $p_{ij}(t)=0$ for all $t>0$.
@@
Let $N\coloneqq\#F$ denote the number of elements in $F$. By enlarging the set $F$ if necessary we may assume that $N$ is an even number. We define the step size $s\coloneqq\frac{t_{0}}{4N}$ and the abbreviations $N_{0}\coloneqq 8\delta N$ and $N_{1}\coloneqq 8(1-\delta)N$. Note that $N_{0},N_{1}\in\mathds{N}$, $N_{0}s=\delta 2t_{0}$, $N_{1}s=(1-\delta)2t_{0}$ and $(N_{0}+N_{1})s=2t_{0}$.
@@
By definition of $C_{m}$, $u(m+1,n)\leq u(m,n)$ for all $m,n$ and $u(m,n)=0$ whenever $n\leq m$. Moreover, $v(m,n)<\varepsilon$ whenever $D_{m}$ is disjoint from $F$ and $n\in\{N_{0},\dots,N_{1}\}$. If $M,m\in\mathds{N}$ with $m<M$, $k\in C_{M}$ and $l\in I$ are such that $p_{lk}(ms)>0$, then $l\in C_{M-m}$. In fact, otherwise there is $M^{\prime}\in\mathds{N}$ with $m<M^{\prime}\leq M$ such that $p_{il}((M^{\prime}-m)s)>0$; this would imply that
$p_{ik}(M^{\prime}s)\geq p_{il}\big{(}(M^{\prime}-m)s\big{)}p_{lk}(ms)>0$
and thus $k\not\in C_{M}$. From this, we now conclude the following two assertions: First,
@@
Recall that at most $N$ of the disjoint sets $D_{m}$ can have a non-empty intersection with $F$. As the set $\{N_{0},\dots,4N-N_{0}\}$ has $4N-2N_{0}+1\geq(4-16\delta)N=3N$ elements, we find $\mathscr{M}\subseteq\{N_{0},\dots,4N-N_{0}\}$ such that #$\#\mathscr{M}=N$ and $D_{m}\subseteq I\setminus F$ for all $m\in\mathscr{M}$. Summing over $m\in\mathscr{M}$, we obtain from (2.1)
$Nc$ $\leq\sum_{m\in\mathscr{M}}\sum_{n=N_{0}+1}^{N_{1}}v(m,n)=\sum_{n=N_{0}+1}^{N_{1}}\sum_{m\in\mathscr{M}}v(m,n)$
@@
$<\sum_{n=N_{0}+1}^{N_{1}}\varepsilon=(N_{1}-N_{0})\varepsilon=(8N-16\delta N)\varepsilon=7N\varepsilon<Nc,$
@@
Let $p_{ij}$ be a positive matrix semigroup over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t_{0})=0$. Then $p_{ij}(t)=0$ for all $t\geq t_{0}$
@@
In order to obtain Theorem 1.1, it remains to show that $p_{ij}(t_{1})>0$ implies that $p_{ij}(t)>0$ for all $t\geq t_{1}$. If the matrix semigroup is standard, this follows easily from $p_{ij}(t_{1}+s)\geq p_{ij}(t_{1})\big{(}p_{jj}(s/n)\big{)}^{n}$ for all $s>0$ and $n\in\mathds{N}$; in case of Markovian matrices the proof is more complex but still based on the fact that $\lim_{t\to 0}p_{jj}(t)$ always exists (but might be zero), cf. [8, Thm II.1.4]. In the general setting of this article this assertion is surprisingly sophisticated to prove. This is the aim of what follows.
@@
Let $p_{ij}$ be a positive matrix semigroup over a countable index set $I$ and let $j\in I$ and assume that $p_{jj}(t)=0$ for all $t>0$.
@@
The assumption $p_{ij}(t_{1})>0$ for some $t_{1}>0$ in Part (b) of Lemma 2.3 may seem artificial at first glance. However, the example
@@
of a constant matrix semigroup shows that this assumption is indeed necessary. The next theorem contains the article’s main result.
@@
Let $p_{ij}$ be a positive matrix semigroup over a countable index set $I$. Let $i,j\in I$ and $t_{0}>0$ such that $p_{ij}(t_{0})=0$. Then $p_{ij}(t)=0$ for all $t>0$.
@@
and $p_{kj}(u-s)>0$ that $p_{kk}(s)=0$. Now Theorem 2.1 implies that $p_{kk}(t)=0$ for all $t>0$. Therefore, $p_{kl}(t)=0$ for all $t>0$ and $l\in I$ by Part (b) of Lemma 2.3. In particular, $p_{kj}$ is identically zero in contradiction to our assumption.
@@
3. Operator Semigroups on Sequence Spaces
@@
In what follows, we prove a version of Theorem 2.4 for one-parameter semigroups on sequence spaces. A family of linear operators $(T_{t})_{t\in(0,\infty)}$ on a vector space is called a (one-parameter operator) semigroup if $T_{t+s}=T_{t}T_{s}$ for all $t,s>0$.
@@
Semigroups appear naturally to describe solutions of linear autonomous evolution equations. In many applications the underlying Banach space is a function space and thus exhibits some kind of order structure. If, in such a situation, a positive initial value of the evolution equation leads to a positive solution, one speaks of a positive semigroup. We refer to the recent monograph [5] for an introduction to the theory of positive semigroups.
In order to make the notion of “sequence spaces” precise, we start with the vector space of all real sequences $\ell^{0}$, which is a vector lattice (or Riesz space) with respect to the component-wise ordering. We refer to [15] and [19] for an introduction to the theory of Riesz spaces and Banach lattices, respectively. For $k\in\mathds{N}$, we write
$e^{k}\coloneqq\big{(}\mathds{1}_{\{k\}}(n)\big{)}_{n\in\mathds{N}}$
for the $k$-th unit vector in $\ell^{0}$. Now we call a sublattice of $\ell^{0}$, i.e. a linear subspace which is closed under lattice operations, a sequence space if it contains all unit vectors $e^{k}$, $k\in\mathds{N}$. All classical “sequence spaces”, such as $\ell^{p}$ ($p$-summable sequences for $1\leq p<\infty$), $\ell^{\infty}$ (bounded sequences), $c$ (convergent sequences) or $c_{0}$ (null sequences), are sequence spaces in the sense of our definition. However, we note that $\ell^{0}$ also has many sublattices which are no sequence spaces. For instance, $C(\mathds{R})$ is as a vector lattice isomorphic to
$C(\mathds{Q})\coloneqq\{g\in\ell^{\infty}(\mathds{Q}):\text{ there exists }f\in C(\mathds{R})\text{ such that }f|_{\mathds{Q}}=g\},$
which in turn is a sublattice of $\ell^{0}$.
Let $E$ be a sequence space. We write $E_{+}\coloneqq\{x\in E:x_{n}\geq 0\text{ for all }n\in\mathds{N}\}$ for its positive cone and recall that a linear operator $T\colon E\to E$ is called positive if $TE_{+}\subseteq E_{+}$. A semigroup is said to be positive if it consists of positive operators. A sequence $(x^{k})\subseteq E$ is said to be order convergent to $x\in E$ if it is order bounded and $x^{k}_{n}\to x_{n}$ for all $n\in\mathds{N}$ as $k\to\infty$. We note that for every $x\in E$, $\sum_{n=1}^{N}x_{n}e^{n}$ is order bounded by $\lvert x\rvert$ and thus order convergent to $x$ as $N\to\infty$. A linear operator $T\colon E\to E$ is sequentially order continuous if $(Tx^{k})$ order converges to $Tx$ whenever $(x^{k})$ order converges to $x$ in $E$.
Using this terminology, Theorem 2.4 becomes the following.
@@
Let $E$ be a sequence space and $(T_{t})_{t\in(0,\infty)}$ be a positive semigroup on $E$ of sequentially order continuous operators. If $t\mapsto(T_{t}e^{i})_{j}$ is continuous on $(0,\infty)$ for all $i,j\in\mathds{N}$, then for all $i,j\in\mathds{N}$ either $(T_{t}e^{i})_{j}>0$ for all $t>0$ or $(T_{t}e^{i})_{j}=0$ for all $t>0$.
Proof.
Let $t,s>0$ and $i,j\in\mathds{N}$. Since $T_{t}$ and $T_{s}$ are order continuous,
$(T_{t+s}e^{i})_{j}=\biggl{(}T_{s}\sum_{k=1}^{\infty}(T_{t}e^{i})_{k}e^{k}\biggr{)}_{j}=\sum_{k=1}^{\infty}(T_{t}e^{i})_{k}(T_{s}e^{k})_{j}.$
This shows that $p_{ij}(t)=(T_{t}e^{i})_{j}$ defines a positive matrix semigroup and Corollary 3.1 follows from Theorem 2.4. ∎
@@
Although the setting of Corollary 3.1 seems to be more special, Theorem 2.4 can be recovered from Corollary 3.1 as follows: Let $p_{ij}$ be a positive matrix semigroup over the index set $\mathds{N}$. Then
$E\coloneqq\{x\in\ell^{0}:\sum\nolimits_{n\in\mathds{N}}\lvert x_{n}\rvert p_{nm}(t)<\infty\text{ for all }t>0\text{ and }m\in\mathds{N}\}$